OpenStudy (anonymous):

For 2x – 3y = –7, when x = 4, What is the value of y?

hero (hero):

2x – 3y = –7 2(4) - 3y = -7 8 - 3y = -7 8 + 7 = 3y 15 = 3y 15/3 = y 5 = y

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

2(4) - 3y = -7 8 - 3y = -7 - 3y - -15 y = 5

OpenStudy (anonymous):

2x – 3y = –7, when x = 4 3y = 2x + 7 Plug x = 4: 3y = 8 + 7 ->y = 15/3 = 5